OpenStudy (anonymous):

what is the square root of 2a to the 3rd power where a=2?

OpenStudy (anonymous):

\[\sqrt{2a ^{3}} \] where with a=2 would be \[\sqrt{2\times(2^{3})} = \sqrt{2^{4}} = \sqrt{16} = 4\]